What is the pKa of benzoic acid if Ka for benzoic acid 6.3 x 10.5?


1
Expert's answer
2021-08-22T07:29:36-0400

Solution:

pKa = - log(Ka)

pKa = - log(6.3×10-5) = -(-4.20) = 4.20

pKa = 4.20


Answer: The pKa of benzoic acid is 4.20